Calculate (a) Gross National Product at MP from the following data.

ITEMS

 (Rs. in Crores)

i. Net factor income to abroad.

10

ii. Private income

1700

iii.Operating surplus.

300

iv. Corporation tax

150

v. Undistributed profit.

30

vi. M I

500

vii. C F C

100

viii. Personal Taxes

200

ix. Compensation of employee.

1200

x. N I T

250


1
Expert's answer
2021-05-25T17:01:35-0400
"Solution"


Using income method :

"NDP_fc = Compensation \ of\ Employees + Operating \ Surplus + MI = 1200 + 300 + 500 = 2000 \\

GNP_mp = NDP_fc + dep + NIT+\ Net\ Factor\ Income \ to\ Abroad = 2000 + 100 + 250 + (-10)\\

 GNP_mp = Rs\ 2340"